# A numerical approach to the Bell-CHSH inequality in Quantum Field Theory

### v1

@@ Abstract
The Bell-CHSH inequality in the vacuum state of a relativistic scalar quantum field is analyzed. Using Weyl operators built with smeared fields localized in the Rindler wedges, the Bell-CHSH inequality is expressed in terms of the Lorentz invariant inner products of test functions. A numerical framework for these inner products is devised. Causality is also explicitly checked by a numerical evaluation of the Pauli-Jordan function. Violations of the Bell-CHSH inequality are reported for different values of the particle mass parameter.
@@ I Introduction
Nowadays, there is a significant interest in the investigation of the Bell-CHSH inequalities in the high-energy particle physics context [15], enabling to study entanglement at energy scales never explored before. Many experimental tests of the Bell-CHSH inequality in the high-energy regime have been proposed in the last years [16, 17, 18, 19, 20, 21]. Quite recently, the first observation of entanglement in a pair of quarks has been reported, being the highest-energy observation of entanglement to date [23].
@@
Since the pioneering works [24, 25, 26], the study of QFT aspects of the Bell-CHSH has received increasing attention, see [27, 28, 29, 30] and references therein for recent accounts. Facing the Bell-CHSH inequality in QFT forces us to dive deeper into the complex structure of quantum fields and their intrinsic relationship with the Minkowski spacetime. As shown in [24, 25], the Bell-CHSH turns out to be already violated in the vacuum state $|0\rangle$, when the field operators are localized in the Rindler wedges.
@@ II Quantum fields and Weyl operators
$\Delta(f,g)$ $=\int\!d^{2}xd^{2}yf(x)\Delta(x-y)g(y),$ (6)
$H(f,g)$ $=\int\!d^{2}xd^{2}yf(x)H(x-y)g(y),$ (7)
with $\Delta(x-y)$ and $H(x-y)$ given by
$i\Delta(x-y)$ $=\frac{1}{2}\int\!\frac{dk}{2\pi}\left(e^{-ik_{\mu}(x-y)^{\mu}}-e^{+ik_{\mu}(x-y)^{\mu}}\right),$
$H(x-y)$ $=\frac{1}{2}\int\!\frac{dk}{2\pi}\left(e^{-ik_{\mu}(x-y)^{\mu}}+e^{+ik_{\mu}(x-y)^{\mu}}\right).$ (8)
@@
The Pauli-Jordan distribution is Lorentz-invariant and encodes the information of locality and relativistic causality, being vanishing outside of the light cone. Moreover, $\Delta(x)$ and $H(x)$ are respectively odd and even under the change $x\rightarrow-x$. Using the smeared fields, we can write $\left[\phi(f),\phi(g)\right]=i\Delta(f,g)$, allowing us to recast causality in this setting as having $\left[\phi(f),\phi(g)\right]=0$ when the supports of $f$ and $g$ are spacelike separated.
@@ IV Numerical setup
Gaussians are smooth but do not have compact support, being in principle not appropriate to be used as test functions. Nevertheless, this issue can be handled with the help of the Pauli-Jordan function. To illustrate this point, let $(h_{R},h_{L})$ stand for two generic Gaussians located in ${\cal W}_{R}$ and in ${\cal W}_{L}$, respectively. The first observation relies on the possibility of shaping $(h_{R},h_{L})$ by a convenient choice of the widths, in such a way that they resemble very much compactly supported functions, meaning that they decay very fast outsides spacetime regions located in ${\cal W}_{R}$ and ${\cal W}_{L}$. Doing so, one can check if $(h_{R},h_{L})$ can be effectively considered to be spacelike by looking at the Pauli-Jordan function $\Delta(h_{R},h_{L})$ and demanding that.
@@
The steps of the method adopted can be resumed as follows. We randomly choose points $(t_{0},x_{0})$ and $(\tau_{0},y_{0})$ in the upper part of ${\cal W}_{R}$, as well as real positive values for $\alpha$ and $\beta$, sampling these parameters in the range $(0,10)$. We also fix the widths $\delta t=\delta x=\delta\tau=\delta y=0.1$ for definiteness. We now pick up a value for the mass $m$ and evaluate all the Hadamard functions needed for the Bell-CHSH inequality (III). Next, with the same set of parameters, we test all the relevant Pauli-Jordan functions, in order to check if causality is fulfilled, according to the threshold (25). All integrals are estimated numerically using the software Mathematica [^1] We mainly used method DoubleExponentialOscillatory, but we also checked the consistency of the numerical integration by adopting other methods.. Then, we check if for this set of parameters a violation of the Bell-CHSH inequality is found, that is, if $\lvert\langle 0|\mathcal{C}|0\rangle\lvert>2$. We repeat then the procedure $5\times 10^{5}$ times for each mass value [^2] To give an idea about the computing time on a single laptop, each mass value required around 15 days to run the tests..
@@
Table 1: Maximum violation achieved for each mass value. For the corresponding Pauli-Jordan functions, we have always $\Delta\leq 10^{-10}$.
@@ V Conclusions
In this work we developed a numerical setup for studying the Bell-CHSH inequality in the vacuum state of a relativistic scalar massive QFT in $(1+1)$ Minkowski spacetime. We defined a Lorentz-invariant inner product through the Wightman smeared two-point function and introduced the Bell-CHSH inequality using the Weyl operators. The integrals present in the inequality were computed numerically and we devised a random test to search for parameters leading to Bell-CHSH inequality violations for different values of the particle mass, see Table (1). The size of the violation decreases with the increase of the mass, suggesting a deep relationship with the cluster property of massive QFT. The setup is of general applicability, being easily adapted to higher dimensional spacetimes as well as to fermionic theories, and opens up the way to several further studies in this domain.

### v2

# Numerical approach to the Bell-CHSH inequality in Quantum Field Theory

@@ Abstract
The Bell-CHSH (Clauser-Horne-Shimony-Holt) inequality in the vacuum state of a relativistic scalar quantum field is analyzed. Using Weyl operators built with smeared fields localized in the Rindler wedges, the Bell-CHSH inequality is expressed in terms of the Lorentz invariant inner products of test functions. A numerical framework for these inner products is devised. Causality is also explicitly checked by a numerical evaluation of the Pauli-Jordan function. Violations of the Bell-CHSH inequality are reported for different values of the particle mass parameter.
@@ I Introduction
Nowadays, there is a significant interest in the investigation of the Bell-CHSH inequalities in the high-energy particle physics context [15], enabling to study entanglement at energy scales never explored before. Many experimental tests of the Bell-CHSH inequality in the high-energy regime have been proposed in the last years [16, 17, 18, 19, 20, 21, 22]. Quite recently, the first observation of entanglement in a pair of quarks has been reported, being the highest-energy observation of entanglement to date [23].
@@
Since the pioneering works [24, 25, 26], the study of QFT aspects of the Bell-CHSH has received increasing attention, see [27, 28, 29, 30, 31, 32, 33] and references therein for recent accounts. Facing the Bell-CHSH inequality in QFT forces us to dive deeper into the complex structure of quantum fields and their intrinsic relationship with the Minkowski spacetime. As shown in [24, 25], the Bell-CHSH turns out to be already violated in the vacuum state $|0\rangle$, when the field operators are localized in the Rindler wedges.
@@ II Quantum fields and Weyl operators
$\Delta(f,g)$ $=\int\!d^{2}xd^{2}yf(x)\hat{\Delta}(x-y)g(y),$ (6)
$H(f,g)$ $=\int\!d^{2}xd^{2}yf(x)\hat{H}(x-y)g(y),$ (7)
with $\hat{\Delta}(x-y)$ and $\hat{H}(x-y)$ given by
$i\hat{\Delta}(x-y)$ $=\frac{1}{2}\int\!\frac{dk}{2\pi}\left(e^{-ik_{\mu}(x-y)^{\mu}}-e^{+ik_{\mu}(x-y)^{\mu}}\right),$
$\hat{H}(x-y)$ $=\frac{1}{2}\int\!\frac{dk}{2\pi}\left(e^{-ik_{\mu}(x-y)^{\mu}}+e^{+ik_{\mu}(x-y)^{\mu}}\right).$ (8)
@@
The Pauli-Jordan distribution is Lorentz-invariant and encodes the information of locality and relativistic causality, being vanishing outside of the light cone. Moreover, $\hat{\Delta}(x)$ and $\hat{H}(x)$ are respectively odd and even under the change $x\rightarrow-x$. Using the smeared fields, we can write $\left[\phi(f),\phi(g)\right]=i\Delta(f,g)$, allowing us to recast causality in this setting as having $\left[\phi(f),\phi(g)\right]=0$ when the supports of $f$ and $g$ are spacelike separated.
@@ IV Numerical setup
Gaussians are smooth but do not have compact support, being in principle not appropriate to be used as test functions. Nevertheless, this issue can be handled with the help of the Pauli-Jordan function. To illustrate this point, let $(h_{R},h_{L})$ stand for two generic Gaussians located in ${\cal W}_{R}$ and in ${\cal W}_{L}$, respectively. The first observation relies on the possibility of shaping $(h_{R},h_{L})$ by a convenient choice of the widths, in such a way that they resemble very much compactly supported functions, meaning that they decay very fast outside spacetime regions located in ${\cal W}_{R}$ and ${\cal W}_{L}$. Doing so, one can check if $(h_{R},h_{L})$ can be effectively considered to be spacelike by looking at the Pauli-Jordan function $\Delta(h_{R},h_{L})$ and demanding that.
@@
The steps of the method adopted can be resumed as follows. We randomly choose points $(t_{0},x_{0})$ and $(\tau_{0},y_{0})$ in the upper part of ${\cal W}_{R}$, as well as real positive values for $\alpha$ and $\beta$, sampling these parameters in the range $(0,10)$ using the Mathematica command “RandomReal”. We also fix the widths $\delta t=\delta x=\delta\tau=\delta y=0.1$ for definiteness. We now pick up a value for the mass $m$ and evaluate all the Hadamard functions needed for the Bell-CHSH inequality (III). Next, with the same set of parameters, we test all the relevant Pauli-Jordan functions, in order to check if causality is fulfilled, according to the threshold (25). All integrals are estimated numerically using the software Mathematica [^1] We mainly used method DoubleExponentialOscillatory, but we also checked the consistency of the numerical integration by adopting other methods.. Then, we check if for this set of parameters a violation of the Bell-CHSH inequality is found, that is, if $\lvert\langle 0|\mathcal{C}|0\rangle\lvert>2$. We repeat then the procedure $5\times 10^{5}$ times for each mass value [^2] To give an idea about the computing time on a single laptop, each mass value required around 15 days to run the tests..
@@
Table 1: Maximum violation achieved for each mass value. For the corresponding Pauli-Jordan functions, we always have $\Delta\leq 10^{-10}$.
@@ V Conclusions
In this work we developed a numerical setup for studying the Bell-CHSH inequality in the vacuum state of a relativistic scalar massive Quantum Field Theory in $(1+1)$ Minkowski spacetime. We defined a Lorentz-invariant inner product through the Wightman smeared two-point function and introduced the Bell-CHSH inequality by means of Weyl operators. The integrals present in the terms of the inequality were computed numerically and we devised a random test in order to search for parameters leading to Bell-CHSH inequality violations for different values of the particle mass, see Table (1).
Let us end by elaborating on the following topics, all under current investigation:
  • As already mentioned, in the case of a relativistic scalar field, several results have been achieved on the violation of the Bell-CHSH inequality in the vacuum state [24, 25, 30], relying on a set of mathematical tools such as von Neumann algebras, Tomita-Takesaki modular theory and the properties of Weyl operators. Those results are of a theoretical nature, aiming at establishing general features of the Bell-CHSH inequality. From this perspective, the construction of an explicit numerical setup is welcome. To our knowledge, this is the first fully numerical attempt to face the Bell-CHSH inequality for a scalar field, including the handling of the check of causality through the Pauli-Jordan function. At the present stage, besides a few technical points, we do not see major obstacles to generalize the present framework to the case of a scalar field in $1+3$ Minkowski spacetime as well as to the case of spinor fields. A point which deserves a certain care is the fact that, in $1+3$ dimensions, the scalar product, Eq. (9), would read
    $\langle f|g\rangle$ $=$ $\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\frac{1}{2\omega_{k}}f^{*}(\omega_{k},{\vec{k}})g(\omega_{k},{\vec{k}}),$ (28)
    $=$ $\int\!\!\frac{{k^{2}}\sin{\theta}dk\;d\theta\;d\varphi}{2(2\pi)^{3}\omega_{k}}\;f^{*}g$
    where
    $f=f(\sqrt{k^{2}+m^{2}},$ $\;k\sin{\theta}\cos{\varphi},$
    $\,k\sin{\theta}\sin{\varphi},\,k\cos{\theta}),$ (29)
    and similarly for $g$. The above expression exhibits the challenges of the $1+3$ case. Multiple integrals need to be evaluated numerically, a feature which might require a number of additional tools. This issue is under consideration and will be reported as soon as possible.
  • As shown in Table (1), the size of the violation decreases with the increase of the mass $m$. This result is deeply related to the cluster property, a general feature of the correlation functions of Quantum Field Theory. Due to its relevance, it is worth to provide here a more detailed discussion. As far as the Bell-CHSH inequality is concerned, we can distinguish basically two cases. The first one is that faced in the present work. More precisely, the test functions $f$ and $g$ refer to two regions, $O$ and $O^{\prime}$, located, respectively, in the right and left Rindler wedges, $({\cal W}_{R},{\cal W}_{L})$, Eq. (22). Essentially, $O$ and $O^{\prime}$ are identified with the two tiny regions in which $f$ and $g$ are effectively nonvanishing, Eq. (25). In this case $O$ and $O^{\prime}$ are not the causal complement of each other. A spatial separation between them can be easily visualized. In such a situation, the Bell-CHSH is expected to display a decay with the increase of the mass, in agreement with the cluster property, see [24]. This is exactly what has been detected: a clear dependence from the mass parameter $m$. There is, however, a second situation which requires a fully different numerical handling. It is the case in which the two regions $O$ and $O^{\prime}$ coincides with the whole wedges $({\cal W}_{R},{\cal W}_{L})$. This means that the test functions $f$ and $g$ are smooth functions supported in $({\cal W}_{R},{\cal W}_{L})$, which vanish outside of $({\cal W}_{R},{\cal W}_{L})$. An example of such a test functions is given, for instance, by
    $f(t,x)=\left\{\begin{aligned} &e^{-\frac{1}{x^{2}-t^{2}}}\;e^{-\frac{x^{2}}{2}}\quad&x\geq|t|\\ &0\quad&{\rm elsewhere}\end{aligned}\right.$ (30)
    where the exponential factor $e^{-\frac{x^{2}}{2}}$ ensures the needed decaying behavior at $x\rightarrow\infty$. The function $f$ in expression (30) is smooth and defined in the whole right wedge ${\cal W}_{R}$, which vanishes at the boundary, i.e. $t=\pm x$. Analogously, in the left wedge ${\cal W}_{L}$, one might consider
    $g(t,y)=\left\{\begin{aligned} &e^{-\frac{1}{y^{2}-t^{2}}}\;e^{-\frac{y^{2}}{2}}\quad&-y\geq|t|\\ &0\quad&{\rm elsewhere}\end{aligned}\right.$ (31)
    This situation is different from the previous one. Here, we are dealing with two regions, $({\cal W}_{R},{\cal W}_{L})$, which are the causal complement of each other. Moreover, there is no spatial separation between them as they touch at the origin, $x=t=0$. For such a spacetime configuration, the theoretical results [24, 25] predict that the violation of the Bell-CHSH inequality achieves its maximum value, namely $2\sqrt{2}$. Unfortunately, the analytic expression of the Fourier transformation of expressions (30) and (31) is not available. As a consequence, a new different numerical setup would be needed. Here, we limit ourselves to mention that we are trying to figure out if a possible framework based on configuration space and not on momentum space might be viable.
    From this discussion, the various numerical aspects and challenges related to the issue of the cluster property should become more apparent. In particular, the geometrical configuration faced in the present work exhibits a clear connection with the cluster property, as evidenced by the results of Table (1).